# Frequency-Domain Characterization of Load Demand from Electrified Highways

### v1

@@ Abstract
Electrified roadways (ER) equipped with dynamic wireless power transfer (DWPT) capabilities can patently extend the driving range and reduce the battery size of electric vehicles (EVs). However, due to the spatial arrangement of the transmitter coils in the ER, the DWPT load exhibits frequency content that could excite power system frequency dynamics. In this context, this work aims to study the spectrum of DWPT loads under different traffic conditions. We develop statistical models for EVs moving at constant speeds to identify the location and magnitude of DWPT load harmonics. Our analysis reveals that the fundamental frequency is dependent on the ER coil spacing and the average EV speed. In the worst-case yet unlikely scenario that EVs move in a synchronized fashion, the amplitude of harmonics scales with the number of EVs. On the contrary, when EVs move freely, harmonics scale with the square root of the number of EVs. Platoon formations can accentuate harmonics. We also show that for higher-order harmonics, the spectral content around harmonics decreases in magnitude and increases in bandwidth. Despite the simplified models, our analysis offers valuable insights for ER planners and grid operators. Numerical tests using a traffic simulator corroborate some of these insights.
Index Terms:
EV charging, oscillations, power spectra, power system dynamics, statistical analysis.
@@ I Introduction
Electrified roadways (ERs) with DWPT capabilities have been proposed to power EVs during transit. The power can be used directly for vehicle propulsion and to charge a battery if needed. DWPT-enabled ERs are expected to offer multiple benefits, such as extended EV range, reduced battery size, and reduced overall battery charging time [1]. DWPT systems may also contribute to increasing EV adoption along with potential cost benefits [2]. From a power grid perspective, with growing EV penetration, stationary EV charging is expected to increase the load on distribution systems significantly [3]. For instance, reference [4] shows that even if 11% of the total number of heavy-duty EVs in Texas charge simultaneously, they would create significant voltage violations. While stationary EV charging loads are concentrated at specific grid buses, DWPT-enabled ERs can distribute EV load across multiple buses.
@@
The DWPT system relies on an array of transmitters (Tx) embedded in the ER that supply power to receivers (Rx) installed on EVs. Successive Tx coils on the ER are separated by a gap [5]. Owing to the spatial arrangement of the Tx coils, the power consumed by each EV exhibits a time-varying, near-periodic waveform. A segment of an ER draws power from the power grid through a substation. The load at the substation equals the total power drawn by all EVs while driving on this ER segment, plus losses. This work aims at modeling the total DWPT load consumed at this substation in the frequency domain.
@@
Previous works predict the average amount of power consumed by the ER infrastructure [6]. Such DWPT load predictions are crucial for system operators to conduct unit commitment, economic dispatch, and power flow studies, as well as determine market prices [7]. Using similar DWPT load predictions, researchers have also designed battery storage systems and control mechanisms to regulate voltage on the distribution grid feeding the ER [8]. Reference [9] studies the optimal location of DWPT lanes, and [10] investigates the vehicle-to-grid power transfer capabilities during blackouts.
@@
While hourly-averaged load profiles of DWPT loads are essential for grid scheduling, they cannot predict the impact of ERs on power grid frequency dynamics. Because stationary EV loads are approximately constant over finer time scales, they have not been a concern for exciting power system oscillations. On the contrary, dynamic EV loads are time-varying and thus pose the risk of introducing oscillations into the power grid. To model the dynamic impact of a DWPT system, its load demand needs to be studied at the timescale of seconds. By assessing the shape of the DWPT frequency spectrum, grid operators can evaluate whether ERs could introduce frequency oscillations into transmission systems.
@@
Transmission system operators are generally interested in identifying possible sources of sustained low-frequency oscillations, also known as inter-area oscillations [11]. These oscillations generally occur at frequencies below 2 Hz and can lead to widespread blackouts if not properly damped; see the recent 2025 Iberian Peninsula blackout [12]. Such oscillations can be triggered due to malfunctions in switching of power electronic devices [13], [14]; sudden loss of load and generators [15]; as well as pulsating electrical loads [16]. Identifying the sources of these oscillations based on real-time measurements is time-intensive. Therefore, it becomes pertinent for power system operators to pre-identify potential sources of such oscillations. This work aims to develop models that determine the frequency spectrum of the total power consumed by a DWPT system under various traffic conditions. Such models are helpful for grid operators and ER designers to understand the impact of DWPT systems on power system dynamics and develop appropriate solutions.
@@
Studies have attempted to predict the real-time power consumption by an ER under diverse traffic flow conditions [17][18]. However, they have not considered the effect of Tx coil segmentation. Reference [1] incorporated the effect of coil segmentation in determining DWPT load demand under different EV penetration levels. Nonetheless, the DWPT spectrum was computed numerically and averaged over diverse traffic conditions, as the time gap between successive EVs was drawn from a distribution inferred based on year-long traffic flow data. Recent work [19] proposes a discrete-time convolution-based method to ascertain the power consumed by DWPT systems, but the analysis does not consider the stochastic nature of traffic.
@@
    We develop statistical models for the DWPT load consumed at a substation serving an ER segment (Section II). The DWPT load is modeled as a continuous-time signal and analyzed in the time and frequency domains. Based on these models, we identify the location and amplitude of frequency harmonics of the DWPT load and study its total harmonic content (THC), defined as the ratio between the non-DC and DC components. Our analysis establishes scaling laws to inform ER designers and grid operators about the potential impact of DWPT loads on grid frequency dynamics.
@@
    Our findings on the DWPT spectrum are summarized as follows (Sec. IIIIV):
@@
    • Assuming vehicles move at a common and constant speed, the highest THC occurs when EVs move in a synchronized fashion. In this worst-case yet unlikely scenario, both the average (DC) DWPT load and its harmonics scale with the number of EVs $N$.
    • When EVs move asynchronously, harmonics do not disappear, but scale with $\sqrt{N}$.
    • Between the two previous extremes, an intermediate case occurs when EVs form platoons with $Q$ EVs per platoon. In this case, harmonics scale with $\sqrt{QN}$.
    • If EVs move at constant yet unequal speeds, the total DWPT load becomes non-periodic. Its spectrum consists of bells centered around harmonics of the average speed. Moreover, the bells decrease in magnitude and increase in bandwidth for higher harmonics.
@@
    Our findings are validated numerically in two ways:
    • Periodograms of randomly sampled DWPT loads resemble the analytically derived spectra (Section V).
    • DWPT loads generated by the SUMO traffic flow simulator demonstrate that free-moving traffic yields mild THC, whereas slower-speed, heavier traffic could lead to synchronization (Sec. VI).
@@ II Time-Domain Model of DWPT Load Demand
This section models the total power consumed by all EVs powered by a DWPT-enabled ER. Hereafter, we will refer to DWPT-enabled EVs simply as EVs. Most existing DWPT studies propose installing the DWPT technology only on a single lane of a highway [18]. We thus assume that DWPT is offered on a single lane and consider one direction of the ER. We consider an ER segment of length $L$ that is powered by a substation. We are interested in modeling the power demand at that substation in the time and frequency domains.
Figure 1: In a DWPT-enabled ER, transmitters Tx of length $\ell_{T}$ and gaps of length $d$ are arranged periodically, every $D=\ell_{T}+d$ meters. A DWPT-enabled EV draws power when its Rx of length $\ell_{R}$ overlaps with Tx coils.
@@
Effects, such as Rx/Tx misalignment, voltage drops across the distribution system feeding the ER, and losses could be captured by adding noise terms in (3), but are ignored for simplicity. The model in (3) also neglects high-frequency harmonics due to the switching of power converters.
@@
We will often refer to the INDOT ER testbed, a pilot project of an ER under development by the Indiana Department of Transportation (INDOT); see [5]. The rightmost lane out of two lanes is instrumented with DWPT technology for a stretch of $400$ m with $\ell_{T}=3.66$ m and $d=0.91$ m. A prototype DWPT-enabled EV has also been developed, featuring an Rx of length $\ell_{R}=1.8$ m. The speed limit is $v=24.6$ m/s or $55$ mph. This testbed is designed to provide up to $200$ kW to an Rx of length $\ell_{R}=1.8$ m, yielding a spatial power density of up to $\bar{a}=111.11$ kW/m.
@@
We are interested in assessing the frequency spectrum of $p(t)$. A key challenge to doing so is that EV positions, speeds, and densities can be random and time-varying. To this end, we postulate statistical models for $p(t)$ to account for various traffic conditions and derive the corresponding spectra. We study four models or scenarios:
@@ III EVs Moving at Common Constant Speeds
Figure 2: DWPT load $p_{n}(t)$ (top) for an EV with Rx length $\ell_{R}=1.8$ m moving at $v=24.65$ m/s ($55$ mph) on the INDOT ER testbed. Blue and green EV loads occur when the EV controller follows the multiplicative model with $a_{n}=\bar{a}$ and $a_{n}=0.75\bar{a}$, respectively. The red EV load occurs when the EV controller clips the power it draws at $0.75\bar{a}$. Because $\ell_{R}>d$, the load is strictly positive at all times. The FS line spectra of blue (middle) and red (bottom) waveforms are also shown. Both spectra peak at harmonics $mf_{0}$ with $f_{0}=5.38$ Hz.
@@
    Although the locations of harmonics $m\omega_{0}$ depend on the EV speed, the FS coefficients $c_{m}$ depend only on Rx/Tx dimensions. To capture the undesirable frequency content, we define the metric of total harmonic content as the ratio
@@
    For the INDOT testbed, signal $h_{T}(t)$ has $\text{THC}_{h}=26\%$.
@@
III-A Scenario S1): Perfect Time Synchronization

### v2

@@ Abstract
Electrified roadways (ER) equipped with dynamic wireless power transfer (DWPT) capabilities can patently extend the driving range and reduce the battery size of electric vehicles (EVs). However, due to the spatial arrangement of the transmitter coils in the ER, the DWPT load exhibits frequency content that could excite power system frequency dynamics. In this context, this work aims to study the spectrum of DWPT loads under different traffic conditions. Under simplifying assumptions, we develop statistical models to identify the location and relative magnitude of DWPT load harmonics. Our analysis reveals that the fundamental frequency depends on ER coil spacing and average EV speed. In the worst-case yet unlikely scenario that EVs move in a synchronized fashion, the amplitude of harmonics scales with the EV count. On the contrary, when EVs move freely, harmonics scale with the square root of the EV count. Platoon formations can accentuate harmonics. The spectral content around harmonics decreases in magnitude and increases in bandwidth with the harmonic index. The load of a single EV moving at a time-varying speed can be modeled as a frequency-modulated (FM) signal. Despite the simplifying assumptions, the derived models offer valuable insights for ER planners and grid operators. Dynamic simulations of a modified WECC model with DWPT loads synthesized from realistic EV trajectories and ER specifications corroborate some of these insights.
@@ I Introduction
Electrified roadways (ERs) with DWPT capabilities have been proposed to power EVs during transit. The power can be used directly for vehicle propulsion and to charge a battery if needed. DWPT-enabled ERs are expected to offer multiple benefits, such as extended EV range, reduced battery size, and reduced overall battery charging time [14]. DWPT systems may also contribute to increasing EV adoption along with potential cost benefits [13]. Currently, the DWPT technology has been proposed for vehicles moving on fixed routes, such as public transit buses and regional-haul Class 8 trucks, as it simplifies infrastructure planning and utilization [5] [29].
From a power grid perspective, with growing EV penetration, stationary EV charging is expected to increase the load on distribution systems significantly [31]. For instance, reference [10] shows that even if 11% of the total number of heavy-duty EVs in Texas charge simultaneously, they would create significant voltage violations. While stationary EV charging loads are concentrated at specific grid buses, DWPT-enabled ERs can distribute EV load across multiple buses.
@@
The DWPT system relies on an array of transmitters (Tx) embedded in the ER that supply power to receivers (Rx) installed on EVs. A segment of an ER draws power from the power grid through a substation. Depending on the ER segment length and anticipated DWPT traffic, the substation can be at the transmission or distribution voltage levels. The substation may serve a segment of an electrified highway and/or additional loads. In most DWPT implementations, successive Tx coils on the ER are separated by a gap [20][7]. Owing to the spatial arrangement of the Tx coils, the power consumed by each EV exhibits a time-varying, near-periodic waveform. The total DWPT load consumed at the substation equals the sum of the EV loads plus losses. This work aims to model the total DWPT load consumed at this substation in the frequency domain.
@@
Previous work predicts the average power consumption of the ER infrastructure [28]. Such DWPT load predictions are crucial for system operators to conduct unit commitment, economic dispatch, and power flow studies, as well as determine market prices [30]. Using similar DWPT load predictions, researchers have also designed battery storage systems and control mechanisms to regulate voltage on the distribution grid feeding the ER [23]. Reference [18] studies the optimal location of DWPT lanes, and [24] investigates the vehicle-to-grid power transfer capabilities during blackouts.
@@
While hourly-averaged load profiles of DWPT loads are essential for grid scheduling, they cannot predict the impact of ERs on power grid frequency dynamics. Because stationary EV loads are approximately constant over finer time scales, they have not been a concern for exciting power system oscillations. On the contrary, dynamic EV loads are time-varying and thus pose the risk of introducing oscillations into the power grid. To model the dynamic impact of a DWPT system, its load demand must be studied at the second timescale. By assessing the shape of the DWPT frequency spectrum, grid operators can evaluate whether ERs could introduce frequency oscillations into transmission systems.
@@
Transmission system operators are interested in identifying possible sources of sustained low-frequency oscillations, also known as inter-area oscillations [12]. These oscillations generally occur at frequencies below 2 Hz and can lead to widespread blackouts if not properly damped; see the recent 2025 Iberian Peninsula blackout [4]. Such oscillations can be triggered due to malfunctions in switching of power electronic devices [8], [2]; sudden loss of load and generators [17]; as well as pulsating electrical loads [21]. Identifying the sources of these oscillations based on real-time measurements is time-intensive. Therefore, it becomes pertinent for power system operators to pre-identify potential sources of such oscillations. This work aims to develop models that determine the frequency spectrum of the total power consumed by a DWPT system under various traffic conditions. Such models are helpful for grid operators and ER designers to understand the impact of DWPT systems on power system dynamics and develop appropriate solutions.
@@
Studies investigating the effects of DWPT systems on the power grid utilize hourly traffic flow data to estimate minute-scale power demand profiles and study voltage violations across a distribution feeder [22]. At this timescale, the effect of Tx coil segmentation can be safely ignored. Reference [14] incorporated the effect of coil segmentation in determining DWPT load demand under different EV penetration levels. Nonetheless, the DWPT spectrum was computed upon averaging across diverse traffic conditions as EV spacings were drawn from a distribution inferred from year-long traffic flow data. Recent work [25] proposes a discrete-time convolution-based method to simulate DWPT loads at a finer timescale, but does not characterize their frequency-domain characteristics.
@@
    We develop statistical models for the DWPT load consumed at a substation serving an ER segment (Sec. II). The DWPT load is modeled as a continuous-time signal and analyzed in the time and frequency domains. Based on these models, we identify the location and amplitude of frequency harmonics of the DWPT load and study its total harmonic content (THC), defined as the ratio between the non-DC and DC components. Our analysis establishes scaling laws to inform ER designers and grid operators about the potential impact of DWPT loads on grid frequency dynamics.
@@
    Our findings on the DWPT spectrum are summarized as follows (Secs. IIIIV):
@@
    • Assuming EVs move at a common and constant speed, the highest THC occurs when EVs move in a synchronized fashion. In this worst-case yet unlikely scenario, both the average (DC) DWPT load and its harmonics scale with the number of EVs $N$. (Scenario S1)
    • When EVs move asynchronously, harmonics scale with $\sqrt{N}$. (Scenario S2)
    • When EVs form platoons with $Q$ EVs per platoon, harmonics scale with $\sqrt{QN}$. (Scenario S3)
    • If EVs move at constant yet unequal speeds, the DWPT load becomes non-periodic. Its spectrum consists of lobes centered around harmonics associated with the average speed. The lobes decrease in magnitude and increase in bandwidth for higher harmonics. Harmonics scale with $\sqrt{N}$ as in S2. (Scenario S4)
    • If EVs move at time-varying speeds, their individual DWPT loads can be modeled as frequency-modulated (FM) signals. Spectra consist of main lobes at harmonics corresponding to the average speed. Each main lobe is accompanied by multiple sidelobes arising due to deviations of speed from the average. (Scenario S5)
@@
    Our findings are validated numerically as follows:
    • Periodograms of randomly sampled DWPT loads resemble the analytically derived spectra (Sec. V).
    • DWPT loads generated by the open-source traffic-flow simulator SUMO [19] demonstrate that the analyses under S1 and S4 explain the periodograms obtained under free-flow and congested traffic, respectively (Sec. VII).
    • DWPT spectrograms (periodograms plotted across time) demonstrate that free-moving traffic yields mild THC, whereas slower-speed heavier traffic could lead to synchronization reaching THC of 18%; see Sec. VII.
    • Feeding DWPT loads obtained from SUMO to the modified WECC 179-bus system demonstrates that during congestion, the DWPT load could excite inter-area oscillations (Sec. VIII). DWPT loads due to free-flowing traffic are unlikely to induce frequency oscillations. Spreading loads across a power system can exacerbate oscillations.
TABLE I: Nomenclature
Symbol Meaning
$L$ length of ER segment
NN number of EVs
$\ell_{T},\ell_{R}$ length of Tx/RX coils
$d$ spacing between successive Tx coils
$D=\ell_{T}+d$ coil segment length
$K=L/D$ number of coil segments
$a_{n}$ spatial power density absorbed by EVn
$x_{n}$ position of head of Rx of EVn
$O(x_{n})$ Rx/Tx overlap for EVn at position $x_{n}$
$\bar{p}_{n}(x_{n})$ power consumed by EVn at position $x_{n}$
$v_{n}(t)$ speed of EVn
$p_{n}(t)$ power consumed by EVn at time $t$
$\bar{a}$ rated power density of an ER
$p(t)$ total power consumed by a DWPT system
$t_{n}$ timing of EVn
$h_{T}(t)$ train of trapezoidal pulses of period $T$
$\omega_{0}=2\pi/T$ fundamental frequency of $h_{T}(t)$
$c_{m}$ $m$-th Fourier series coefficient of $h_{T}(t)$
THCh,THCp THC of $h_{T}(t)$, $p(t)$
$\tilde{c}_{m}$ $m$-th FS coeffient of $p(t)$
$P(\omega)$ Fourier transform of $p(t)$
$\delta(\omega)$ Dirac delta function
$\mu_{p},R_{p}(\tau)$ mean and autocorrelation of $p(t)$
$S_{p}(\omega)=\mathcal{F}[R_{p}(\tau)]$ Power spectral density of $p(t)$
$\mu_{a},\sigma_{a}^{2}$ mean and variance of $a_{n}$
$Q$ number of EVs per platoon
$\lambda_{a}$ mean vehicle arrival rate
$T_{n}=D/v_{n}$ fundamental period of EVn
$\mu_{v},\sigma_{v}^{2}$ mean and variance of speeds $v_{n}$
$\mu_{\omega},\sigma_{\omega}^{2}$ mean and variance of $\omega_{n}$
$\hat{p}(t)$ DWPT load realization of duration $T_{p}$
$\hat{S}_{p}(\omega)$ periodogram of $\hat{p}(t)$
$\bar{v}_{n}$$\tilde{v}_{n}(t)$ mean EV speed, deviations from mean speed
$\omega_{s}=2\pi/D$ spatial fundamental frequency
@@ II Time-Domain Model of DWPT Load Demand
This section models the total power consumed by all EVs powered by a DWPT-enabled ER. Hereafter, we will refer to DWPT-enabled EVs simply as EVs. We assume the DWPT technology is available on a single lane of the ER as suggested by many existing DWPT studies [3]. Moreover, without loss of generality, we consider one direction of the ER, e.g., northbound alone, rather than both northbound and southbound. We consider an ER segment of length $L$ that is powered by a substation. We are interested in modeling the power demand at that substation in the time and frequency domains.
Figure 1: In a DWPT-enabled ER, transmitters Tx of length $\ell_{T}$ and gaps of length $d$ are arranged periodically, every $D=\ell_{T}+d$ meters. An EV draws power when its Rx of length $\ell_{R}$ overlaps with the Tx coils.
@@
Figure 2: The INDOT ER prototype is powered by a substation that steps down grid voltage to $12.47$ kV. Each feeder is connected to a transformer-rectifier unit that converts power from AC to DC. DC power is then fed to inverters installed at the side of each Tx coil [14]. The Tx coils are designed to operate at around $765$ V and can supply $190$ kW of power.
@@
We will often refer to the INDOT ER testbed, a pilot project of an ER under development by the Indiana Department of Transportation (INDOT). The rightmost lane out of two lanes is instrumented with DWPT technology for a stretch of $400$ m with $\ell_{T}=3.66$ m and $d=0.91$ m. A prototype DWPT-enabled EV has also been developed, featuring an Rx of length $\ell_{R}=1.8$ m. The speed limit is $v=24.6$ m/s or $55$ mph. This testbed is designed to provide up to $200$ kW to an Rx of length $\ell_{R}=1.8$ m, yielding a spatial power density of up to $\bar{a}=111.11$ kW/m. Figure 2 shows a schematic of the components powering this ER [1]. The distribution substation transformer steps the voltage down to 12.47 kV, which is then supplied to the transformer-rectifier (TR) units via feeders. Each TR unit steps down the voltage level to around 800-900 V as well as converting power from AC to DC to reduce losses. Tx coils are supplied with AC power via dedicated inverters, one per Tx coil. The losses include ohmic losses along the feeder, leakage in the Rx/Tx electromagnetic coupling, and inefficiencies in DC/AC inverters. The INDOT ER pilot has achieved an 85% Tx-to-Rx power transfer efficiency [20]. Note that this is only one possible DWPT architecture, as there are no standardized designs yet.
@@
Effects, such as Rx/Tx misalignment, voltage drops across the distribution system feeding the ER, and losses, could be captured by adding noise terms in (3) or (5), but are ignored for simplicity. Inverters and TR units operate at a switching frequency of 85 kHz, which introduces high-frequency content to $p_{n}(t)$. Because the transmission network is a natural low-pass filter, such content would not excite grid frequency dynamics and can be ignored. Interested readers can refer to [20] for precise DWPT voltage and power waveforms.
We aim to assess the frequency spectrum of $p(t)$. An important summarization of the DWPT load spectrum is the ratio of the undesirable frequency content of $p(t)$ to its DC component. We term this metric total harmonic content and define it formally as
$\text{THC}_{p}=\frac{\sqrt{\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}\left(p(t)-\bar{p}\right)^{2}dt}}{\bar{p}}\times 100\%,$ (6)
where $\bar{p}$ is the time-averaged DWPT load:
$\bar{p}=\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}p(t)dt.$
The DC component is the actual power consumed by the DWPT load. The non-DC component comprises fluctuations in $p(t)$ about its average value. How are such fluctuations distributed across frequencies? How does THCp change with traffic conditions? Depending on their frequency content and THC, DWPT loads could excite grid frequency dynamics.
A key challenge in estimating the frequency spectrum of the DWPT load is that EV positions, speeds, and densities are random and time-varying. To this end, we postulate statistical models for $p(t)$ to account for various traffic conditions and derive the corresponding spectra. We derive the frequency spectrum and THC of the total DWPT load under five models or scenarios:
@@ 4. S4)
  5. S5)
    Model the spectrum of a single EV moving with time-varying speed.
@@ III EVs Moving at Common Constant Speeds
Figure 3: Top: The DWPT load $p_{n}(t)$ for an EV with Rx length $\ell_{R}=1.8$ m moving at $v=24.65$ m/s ($55$ mph) on the INDOT ER testbed. The blue and green load waveforms occur when the EV controller follows the multiplicative model with $a_{n}=\bar{a}$ and $a_{n}=0.75\bar{a}$, respectively. The red load waveform occurs when the EV controller clips the drawn power at $0.63\bar{a}$. The maximum power of the clipped DWPT load is chosen such that the DC component of the red and green waveforms is the same. Additionally, since $\ell_{R}>d$, the clipped DWPT load is strictly positive at all times. Bottom: The FS line spectrum of the red and green load waveform shown on the top panel ($c_{0}=0.6$ p.u.). The spectrum peaks at harmonics $mf_{0}$ with $f_{0}=5.38$ Hz. Observe that the clipped DWPT load introduces harmonics of lower amplitude compared to the load generated by the multiplicative model (see [11] for more details on the comparison between different DWPT load models).
@@
    Although the locations of harmonics $m\omega_{0}$ depend on the EV speed, the FS coefficients $c_{m}$ depend only on Rx/Tx dimensions. Applying the definition of THC to the periodic load $p_{n}(t)$ in (12) leads to the alternative THC expression
@@
    For the INDOT testbed, the signal $h_{T}(t)$ has $\text{THC}_{h}=26\%$. The THC should not be confused with the total harmonic distortion (THD), which measures the relative power contained in the non-fundamental frequencies of an AC electric signal relative to its fundamental frequency.
@@
III-A Scenario S1: Perfect Time Synchronization